# Non-Gaussian Photon Correlations in Weakly Coupled Atomic Ensembles

### v1

@@ Abstract
We develop a scattering theory formalism and use it to predict that a resonantly driven atomic ensemble weakly coupled to an optical mode can generate light with non-Gaussian correlations. Our approach—based on a perturbative diagrammatic expansion of multi-photon interactions—shows that photon-photon interaction mediated by the emitters causes the transmitted light to have a non-vanishing connected third-order correlation function $g_{c}^{(3)}$. We explain the temporal pattern of $g_{c}^{(3)}$ using the interaction processes in our diagrammatic expansion. A quantitative comparison with cascaded master equation simulations for small ensembles with optical depth $\mathrm{OD}\leq 2$ confirms that the perturbative results remain accurate across experimentally relevant optical depths and for drive strengths large enough to make the predicted non-Gaussian signatures detectable. We anticipate that state-of-the-art nanofibre-coupled atomic ensembles can experimentally demonstrate our predictions.
@@
Introduction.—Photon-photon interactions in optically driven atomic ensembles can alter the quantum state of an incident laser field. Here, the interplay of nonlinear interactions, due to the two-level atoms in the ensemble, and the elastic scattering of laser light out of the ensemble allows manipulating the quantum nature of light [1, 2]. For example, photons can become strongly correlated, modifying both intensity correlations [3, 4, 5] and electric field fluctuations [6, 7]. Such works have largely focused on quantifying two-body photon correlations. On the other hand, three-photon correlations, including non-Gaussian correlations, have been measured in other quantum optics systems [8, 9, 10, 11, 12], and very recently an atomic ensemble was also used to demonstrate the generation of non-Gaussian light by observing violations of the Siegert relation [13]. Here, non-Gaussian refers to light whose observables or correlators cannot be described by the mean-values and covariance of its electric field quadratures [14].
@@
Despite experimental advances, a tractable theoretical framework for describing three-photon dynamics or the presence of non-Gaussian correlations in large atomic ensembles remains elusive. Dilute ensembles are often modeled using the Maxwell-Bloch equations [15, 16, 17, 18], which, under a mean-field approximation, provide a minimal semiclassical description incorporating unidirectional coupling to a one-dimensional continuum of photon modes. To capture quantum correlations beyond the mean-field level, one can employ a complete cascaded master equation [19, 20]. However, for large ensembles, the exponential growth of the Hilbert space with system size renders the direct computation of higher-order correlation functions, such as $g^{(3)}$, challenging, even if employing cumulant expansion [21, 22] or matrix-product operator [23, 24] methods.
@@
In this work, we propose an analytic and diagrammatic framework to calculate up to three-photon transport through the atomic ensemble. Using the Maxwell-Bloch approach, the ensemble can be considered to be an array of atoms unidirectionally coupled to an optical nanofibre as shown in Fig. 1. The coupling efficiency between each atom and waveguide is weak $\beta\ll 1$ while the number of atoms $M\gg 1$, leading to significant optical depth $\mathrm{OD}=4\beta M$. Our diagrammatic approach takes advantage of the weak coupling and uses $\beta$ as an expansion parameter, which enables us to analytically compute leading-order contributions to the three-photon-correlated component of the wavefunction and the connected third-order correlation function $g_{c}^{(3)}$. We confirm the presence of light with significant non-Gaussian correlations ($|g_{c}^{(3)}|>0.1$) for weakly driven atomic ensembles of moderate optical depth $\mathrm{OD}\sim 2$.
@@
Model.—The model for our system is depicted in Fig. 1, where $M$ atoms are unidirectionally coupled to a propagating channel in a waveguide. Here we have illustrated the example of a nanofibre geometry where atoms are trapped near the fibre and interact via its evanescent field. Each atom couples to the propagating channel at rate $\Gamma=\beta\Gamma_{\rm tot}$ and to its individual loss channel at rate $\gamma=(1-\beta)\Gamma_{\rm tot}$, where $\beta$ quantifies the coupling efficiency and $\Gamma_{\rm tot}$ is the overall atomic decay rate. The emitters are driven by a coherent on-resonance driving field from one end of the waveguide. The emitter spacing exceeds one resonant wavelength of the two-level atoms which ensures negligible collective emission outside the waveguide. Typical experimental values for atomic ensembles coupled to nanofibres yield $\beta\sim 0.01$ [3]. Despite the small individual coupling strength, the OD of the entire atomic array becomes significant $\mathrm{OD}=\mathcal{O}(1)$, where $\mathcal{O}(\cdot)$ indicates big-O notation. This enables the collective nonlinear effects and strong dissipation to substantially modulate photon transport properties, which, as we show, results in strongly correlated photons and non-Gaussian output states of light.
@@
$+$ $\delta(x)\left[\hat{\sigma}_{m}^{+}(\sqrt{\Gamma}\hat{a}(x)+\sqrt{\gamma}\hat{b}_{m}(x))+h.c.\right]\Biggr{\}},$
@@
The $n$-photon scattering in the even subspace can be fully described by the $S$-matrix derived from Bethe’s Ansatz [25, 26]. The full $S$-matrix element encodes all the information of free propagation of photons, photon-atom and photon-photon interactions. We refer to free propagation of photons and individual elastic photon-atom scattering as disconnected processes because they factorize into independent terms in the $S$-matrix. The connected part of $S$-matrix is defined by subtracting all the products of the disconnected processes from the full $S$-matrix, enabling us to extract the genuine effects of $n$-body photon-photon interactions [27]. The expression of the full and connected two- and three-photon $S$-matrix is given in the accompanying work [28].
@@
The local scattering event at the $m$th atom can be calculated by the following steps: (1) transform the incoming state from the propagating channel into even/odd channel picture (2) multiply the term with $n$ even photons by the $n$-photon $S$-matrix (3) transform back to propagating/loss channel picture. When the photon remains in the waveguide, elastic single-photon scattering reduces to multiplication by the $k$-space transmission coefficient $t_{k}=1-i\beta\Gamma_{\mathrm{tot}}/(k+i\Gamma_{\mathrm{tot}}/2)\,$, while two- and three-photon interaction events are captured by the multiplications of their respective connected scattering matrices $\beta^{2}\hat{S}^{C}_{p_{1}p_{2},k_{1}k_{2}}$ and $\beta^{3}\hat{S}^{C}_{p_{1}p_{2}p_{3},k_{1}k_{2}k_{3}}$. When a photon is lost by elastic scattering out of the waveguide, the incoming state is multiplied by a reflection coefficient $r_{k}=-\sqrt{\beta(1-\beta)}i\Gamma_{\rm tot}/(k+i\Gamma_{\rm tot}/2)\,$. Alternatively, if loss occurs immediately after a two- or three-photon interaction, the incoming state is multiplied by $\beta^{3/2}\sqrt{(1-\beta)}\hat{S}^{C}_{p_{1}p_{2},k_{1}k_{2}}$ or $\beta^{5/2}\sqrt{(1-\beta)}\hat{S}^{C}_{p_{1}p_{2}p_{3},k_{1}k_{2}k_{3}}$ respectively.
@@
Expressing the scattering in terms of the connected part of the $S$-matrix clarifies the way in which genuine photon–photon interactions scale with the various powers of coupling strength $\beta$. Since $\beta\ll 1$, this scaling behavior motivates us to perturbatively treat the effect of photon-photon interactions in terms of the order in $\beta$. If a three-photon transport process has no photon-photon interaction, the outgoing state is simply the incoming state of the array multiplied by $t_{0}^{3M}$. Transport processes generating three-photon correlations are the main interest of this work. In Fig. 2, we sketch the interaction processes where three-photon trajectories are connected by interaction lines. These diagrams are drawn according to the following rules: (1) Each horizontal line indicates the trajectory of each photon being transported through the array from the left to the right. The individual elastic atom-photon scatterings are not explicitly indicated on the line. (2) Each black circular dot on the photon trajectory represents an atomic site where the photon interacts with other photons. The dots on the same column corresponds to the same atomic site. (3) A $n$-photon interaction is represented by a wavy line connecting $n$ dots. We note that black dots can be any atom in the array, therefore each diagram in Fig. 2 represents all possible transport processes with the interactions happening on different atomic sites. For example, Fig. 2(a) shows three photons interacting via a single atom. This can occur at any atom within the ensemble and therefore has a combinatorial factor of $M$. Photon-photon interactions at the start of the ensemble are more likely as each elastic scattering before the photon-photon interaction multiplies the photon wavefunction by $t_{0}$, exponentially decreasing its amplitude along the array.
@@
where after $j$ individual single-photon scatterings, the first two-photon interaction occurs at site $j+1$; the middle photon with momentum $q_{2}$ then undergoes $m$ additional individual single-photon scatterings before the second two-photon interaction at site $j+m+2$, followed by the remaining Mjm2M-j-m-2 individual single-photon scatterings. Due to the bosonic symmetry, there are six 4-vertex diagrams corresponding to different permutations of the outgoing momenta $p_{1},p_{2},p_{3}$ at the right end of the diagram. To estimate the order of the amplitude of each diagram for perturbative expansion, we need to consider (1) the weight of the coupling constant $\beta$, and (2) the combinatorics of the process due to interactions potentially occurring at any atom in the ensemble. For (1), we follow how photon-creation operators transform under Weyl transformation for nonlinear interactions. Each transmitted photon picks up a factor $\sqrt{\beta}$ when moving from the propagating to the even channel and another $\sqrt{\beta}$ when transforming back to the propagating channel, giving an overall factor $\beta$. This means the wavefunction amplitude of a single two-photon interaction scales as $\beta^{2}$, while a single three-photon interaction scales as $\beta^{3}$. For a lost photon, the second factor becomes $\sqrt{1-\beta}$, so the vertex contributes $\sqrt{\beta(1-\beta)}\approx\sqrt{\beta}$. For (2), the combinatorics of a diagram with $n$ photon-photon interactions is $\binom{M}{n}$. For larger OD $\beta M=\mathcal{O}(1)$, the magnitude is estimated by multiplying $\beta$ and the combinatorial factor. For example, the order of Fig. 2b is $\binom{M}{2}\beta^{4}\approx M^{2}\beta^{4}=\mathcal{O}(\beta^{2})$. For the order estimation in small or intermediate OD, because the power of $\beta$ is fixed for each diagram, its magnitude grows monotonically with $M$, therefore the leading diagrams at small or intermediate OD are a subset of those at large OD. Hence, perturbative results derived for the large-OD regime remain valid, and even gain accuracy, when applied to lower-OD systems.
From the Eq. 3 and  4, we see that at large OD, $\hat{T}^{3v}$ and $\hat{T}^{4v}$ are both at $\mathcal{O}(\beta^{2})$, which is the leading-order contribution among all possible three-photon-connected diagrams. The subleading order $\mathcal{O}(\beta^{3})$ three-photon connected diagrams includes two types of diagrams: (1) the diagrams with one three-photon interaction and one two-photon interaction, and (2) the diagrams with three two-photon interactions. One representative of each type is shown in Figs. 2c and  2d. These diagram are negligible at $\beta\approx 1\%$, but are significant at large OD with $\beta\approx 5\%$. As the diagrams feature a loop structure, we refer to these as loop-order diagrams and their explicit calculation is presented in our accompanying work [28].
@@
Connected third-order correlation function.—Having identified the leading-order contributions to connected three-photon scattering, we now translate this into an experimentally accessible quantity that indicates the non-Gaussianity of the outgoing light. The connected third-order correlation function is defined as
@@
where $g^{(2)}(x_{1},x_{2})$ and $g^{(3)}(x_{1},x_{2},x_{3})$ are the normalized second- and third-order correlation functions, respectively. For non-vanishing power, $g_{c}^{(3)}(x_{1},x_{2},x_{3})$ vanishes when either one photon is far away from the other two, or when the state of the light is Gaussian [29, 30, 8]. In the weak driving condition $\mathcal{O}(P_{\rm in}/\Gamma_{\rm tot})=\mathcal{O}(\beta)$  [^1] Stricly, we require $\mathcal{O}(P_{\rm in}/\Gamma_{\rm tot})=\mathcal{O}(\beta^{2})$, but it is shown in the accompanying work [28] that this condition can be loosened to be $\mathcal{O}(P_{\rm in}/\Gamma_{\rm tot})=\mathcal{O}(\beta)$., $g^{(n)}(x_{1},\dots,x_{n})=\expectationvalue{:\Pi_{j=1}^{n}\hat{a}^{\dagger}(x_{j})\hat{a}(x_{j}):}/\Pi_{k=1}^{n}\expectationvalue{\hat{a}^{\dagger}(x_{k})\hat{a}(x_{k})}$ for $n=2,3$ can be approximated as:
@@
where $\psi_{2}(x_{1},x_{2})=t_{0}^{2M}+\phi_{2}(x_{1},x_{2})$ with $\phi_{2}(x_{i},x_{j})$ is the entangled wavefunction generated by all possible two-photon interactions. The output power $\expectationvalue{\hat{a}^{\dagger}(x)\hat{a}(x)}\approx P_{\rm in}t_{0}^{2M}$, $\psi_{3}(x_{1},x_{2},x_{3})=t_{0}^{3M}+t_{0}^{M}\sum_{1\leq i<j\leq 3}\,\phi_{2}(x_{i},x_{j})+\phi_{3}(x_{1},x_{2},x_{3})$, where $\phi_{3}$ encodes the correlation induced by leading- and subleading-order three-photon interactions depicted in Fig. 2. We evaluate $\phi_{2}$ and $\phi_{3}$ analytically, but for brevity here, their expressions are given in [28].

### v2

@@ Abstract
We develop a scattering theory formalism and use it to predict that a resonantly driven atomic ensemble weakly coupled to an optical mode can generate light with non-Gaussian correlations. Our approach—based on a perturbative diagrammatic expansion of multi-photon interactions—shows that photon-photon interaction mediated by the emitters causes the transmitted light to have a connected third-order correlation function $g_{c}^{(3)}$ revealing its non-Gaussianity. We explain the temporal pattern of $g_{c}^{(3)}$ using the interaction processes in our diagrammatic expansion. A quantitative comparison with cascaded master equation simulations for small ensembles with optical depth $\mathrm{OD}\leq 2$ confirms that the perturbative results remain accurate across experimentally relevant optical depths and for drive strengths large enough to make the predicted non-Gaussian signatures detectable. We anticipate that state-of-the-art nanofibre-coupled atomic ensembles can experimentally demonstrate our predictions.
@@
Introduction.—Photon-photon interactions in optically driven atomic ensembles can alter the quantum state of an incident laser field. Here, the interplay of nonlinear interactions, due to the two-level atoms in the ensemble, and the elastic scattering of laser light out of the ensemble allows manipulating the quantum nature of light [Mahmoodian2018, RevModPhys.95.015002]. For example, photons can become strongly correlated, modifying both intensity correlations [Prasad2020, Cordier2023, Ferioli2025PRLa] and electric field fluctuations [Mahmoodian2021, Lu1998PRL]. Such works have largely focused on quantifying two-body photon correlations. On the other hand, three-photon correlations, including non-Gaussian intensity and field correlations, have been measured in other quantum optics systems [Stiesdal2018, Liang2018Science, Ornelas-Huerta2021PRL, Tomm2023NPHYS, Das2025arXiv], and very recently an experiment using an atomic ensemble claimed to observe non-Gaussian light via the violation of the Siegert relation [Ferioli2024]. Here, We use the term “non-Gaussian” in two distinct senses. By non-Gaussian field correlations we mean that the output state is non-Gaussian in phase space (equivalently, its Wigner function is not Gaussian), in which case higher-order field moments are not fixed by first and second moments. By non-Gaussian intensity correlations we mean that normally ordered intensity cumulants beyond second order are nonzero, e.g., $g_{c}^{(3)}\neq 0$. Importantly, a Gaussian field state can still display non-Gaussian intensity correlations (thermal light is a standard example).
@@
Despite experimental advances, a tractable theoretical framework for describing three-photon dynamics or the presence of non-Gaussian correlations in large atomic ensembles remains elusive. Dilute ensembles are often modeled using the Maxwell-Bloch equations [McCall1969PR, Hammerer2010RMP, Goncalves2025PRXQ, Kusmierek2024arXiv], which, under a mean-field approximation, provide a minimal semiclassical description incorporating unidirectional coupling to a one-dimensional continuum of photon modes. To capture quantum correlations beyond the mean-field level, one can employ a complete cascaded master equation [Gardiner1993PRL, Carmichael1993PRL]. However, for large ensembles, the exponential growth of the Hilbert space with system size renders the direct computation of higher-order correlation functions, such as $g^{(3)}$, challenging, even if employing cumulant expansion [Mahmoodian2023, Plankensteiner2022Quantum] or matrix-product operator [manzoni2017NCOMM, Mahmoodian2020PRX] methods.
@@
In this work, we propose an analytic and diagrammatic framework to calculate up to three-photon transport through the atomic ensemble. Using the Maxwell-Bloch approach, the ensemble can be considered to be an array of atoms unidirectionally coupled to an optical nanofibre as shown in Fig. 1. The coupling efficiency between each atom and waveguide is weak $\beta\ll 1$ while the number of atoms $M\gg 1$, leading to significant optical depth $\mathrm{OD}=4\beta M$. Our diagrammatic approach takes advantage of the weak coupling and uses $\beta$ as an expansion parameter, enabling us to analytically compute leading-order contributions to the three-photon-correlated component of the wavefunction and the connected third-order correlation function $g_{c}^{(3)}$. So far, our perturbative approach is the only way to analytically compute three-photon interactions, which is the cause of non-Gaussian field correlation, in large atomic ensembles. To show that the outgoing state after transport has non-Gaussian field correlation, we demonstrate that $g_{c}^{(3)}$ calculated from our approach disagrees with $\widetilde{g}_{c}^{(3)}$ calculated by Isserlis’s theorem based on the hypothesis that the state is a Gaussian field.
@@
Model.—The model for our system is depicted in Fig. 1, where $M$ atoms are unidirectionally coupled to a propagating channel in a waveguide. Here we have illustrated the example of a nanofibre geometry where atoms are trapped near the fibre and interact via its evanescent field. Each atom couples to the propagating channel at rate $\Gamma=\beta\Gamma_{\rm tot}$ and to its individual loss channel at rate $\gamma=(1-\beta)\Gamma_{\rm tot}$, where $\beta$ quantifies the coupling efficiency and $\Gamma_{\rm tot}$ is the overall atomic decay rate. The emitters are driven by a resonant coherent driving field from one end of the waveguide. The emitter spacing exceeds one resonant wavelength of the two-level atoms which ensures negligible collective emission outside the waveguide. Typical experimental values for atomic ensembles coupled to nanofibres yield $\beta\sim 0.01$ [Prasad2020]. Despite the small individual coupling strength, the OD of the entire atomic array becomes significant $\mathrm{OD}=\mathcal{O}(1)$, where $\mathcal{O}(\cdot)$ indicates big-O notation. This enables the collective nonlinear effects and strong dissipation to substantially modulate photon transport properties, which, as we show, results in strongly correlated photons and non-Gaussian output states of light.
@@
$+$ $\delta(x)\left[\hat{\sigma}_{m}^{+}(\sqrt{\Gamma}\hat{a}(x)+\sqrt{\gamma}\hat{b}_{m}(x))+{\rm h.c.}\right]\Biggr\},$
@@
The $n$-photon scattering in the even subspace can be fully described by the $S$-matrix derived from Bethe’s Ansatz [Yuds1985, Rupasov1984JETP]. The full $S$-matrix element encodes all the information of free propagation of photons, photon-atom and photon-photon interactions. We refer to free propagation of photons and individual elastic photon-atom scattering as disconnected processes because they factorize into independent terms in the $S$-matrix. The connected part of $S$-matrix is defined by subtracting all the products of the disconnected processes from the full $S$-matrix, enabling us to extract the genuine effects of $n$-body photon-photon interactions [WeinbergQFT]. The expression of the full and connected two- and three-photon $S$-matrix is given in the accompanying work [WDM2025].
@@
The local scattering event at the $m$th atom can be calculated by the following steps: (1) transform the incoming state from the propagating channel into even/odd channel picture (2) multiply the term with $n$ even photons by the $n$-photon $S$-matrix (3) transform back to propagating/loss channel picture. When the photon remains in the waveguide, elastic single-photon scattering reduces to multiplication by the $k$-space transmission coefficient $t_{k}=1-i\beta\Gamma_{\mathrm{tot}}/(k+i\Gamma_{\mathrm{tot}}/2)\,$, [Mahmoodian2018, Yudson2008] while two- and three-photon interaction events are captured by the multiplications of their respective connected scattering matrices $\beta^{2}\hat{S}^{C}_{p_{1}p_{2},k_{1}k_{2}}$ and $\beta^{3}\hat{S}^{C}_{p_{1}p_{2}p_{3},k_{1}k_{2}k_{3}}$. When a photon is lost by elastic scattering out of the waveguide, the incoming state is multiplied by a reflection coefficient $r_{k}=-\sqrt{\beta(1-\beta)}i\Gamma_{\rm tot}/(k+i\Gamma_{\rm tot}/2)\,$. Alternatively, if loss occurs immediately after a two- or three-photon interaction, the incoming state is multiplied by $\beta^{3/2}\sqrt{(1-\beta)}\hat{S}^{C}_{p_{1}p_{2},k_{1}k_{2}}$ or $\beta^{5/2}\sqrt{(1-\beta)}\hat{S}^{C}_{p_{1}p_{2}p_{3},k_{1}k_{2}k_{3}}$ respectively.
@@
Expressing the scattering in terms of the connected part of the $S$-matrix clarifies the way in which genuine photon–photon interactions scale with the various powers of coupling strength $\beta$. Since $\beta\ll 1$, this scaling behavior motivates us to perturbatively treat the effect of photon-photon interactions in terms of the order in $\beta$. If a three-photon transport process has no photon-photon interaction, the outgoing state is simply the incoming state of the array multiplied by $t_{0}^{3M}$. Transport processes generating three-photon correlations are the main interest of this work. In Fig. 2, we sketch the interaction processes where three-photon trajectories are connected by interaction lines. These diagrams are drawn according to the following rules: (1) Each horizontal line indicates the trajectory of each photon being transported through the array from the left to the right. The individual elastic atom-photon scatterings are not explicitly indicated on the line. (2) Each black circular dot on the photon trajectory represents an atomic site where the photon interacts with other photons. The dots on the same column corresponds to the same atomic site. (3) A $n$-photon interaction is represented by a wavy line connecting $n$ dots. We note that black dots can be any atom in the array, therefore each diagram in Fig. 2 represents all possible transport processes with the interactions happening on different atomic sites. For example, Fig. 2(a) shows three photons interacting via a single atom. This can occur at any atom within the ensemble and therefore the amplitude of this type of processes has a combinatorial factor of $M$. Photon-photon interactions at the start of the ensemble are more likely as on-resonant photon transports are lossy. Each elastic scattering before the photon-photon interaction multiplies the photon wavefunction by $t_{0}$, exponentially decreasing its amplitude along the array.
@@
where after $j$ individual single-photon scatterings, the first two-photon interaction occurs at site $j+1$; the middle photon with momentum $q_{2}$ then undergoes $m$ additional individual single-photon scatterings before the second two-photon interaction at site $j+m+2$, followed by the remaining Mjm2M-j-m-2 individual single-photon scatterings. Due to the bosonic symmetry, there are six 4-vertex diagrams corresponding to different permutations of the outgoing momenta $p_{1},p_{2},p_{3}$ at the right end of the diagram. To estimate the order of the amplitude of each diagram for perturbative expansion, we need to consider (i) the weight of the coupling constant $\beta$, and (ii) the combinatorics of the process due to interactions potentially occurring at any atom in the ensemble. For (i), we follow how photon-creation operators transform under Weyl transformation for nonlinear interactions. Each transmitted photon picks up a factor $\sqrt{\beta}$ when moving from the propagating to the even channel and another $\sqrt{\beta}$ when transforming back to the propagating channel, giving an overall factor $\beta$. This means the wavefunction amplitude of a single two-photon interaction scales as $\beta^{2}$, while a single three-photon interaction scales as $\beta^{3}$. For a lost photon, the second factor becomes $\sqrt{1-\beta}$, so the vertex contributes $\sqrt{\beta(1-\beta)}\approx\sqrt{\beta}$. For (ii), the combinatorics of a diagram with $n$ photon-photon interactions is $\binom{M}{n}$. For larger OD $\beta M=\mathcal{O}(1)$, the order of magnitude of a diagram is estimated by multiplying $\beta$ and the combinatorial factor. For example, the order of Fig. 2b is $\binom{M}{2}\beta^{4}\approx M^{2}\beta^{4}=\mathcal{O}(\beta^{2})$. For the order estimation in small or intermediate OD, because the power of $\beta$ is fixed for each diagram, the amplitude of the diagram grows monotonically with $M$, therefore the leading diagrams at small or intermediate OD are a subset of those at large OD. Hence, perturbative results derived for the large-OD regime remain valid, and even gain higher accuracy, when applied to lower-OD systems.
From the Eq. 3 and  4, we see that at large OD, $\hat{T}^{3v}$ and $\hat{T}^{4v}$ are both at $\mathcal{O}(\beta^{2})$, which is the leading-order contribution among all possible three-photon-connected diagrams. The subleading order $\mathcal{O}(\beta^{3})$ three-photon connected diagrams includes two types of diagrams: (1) the diagrams with one three-photon interaction and one two-photon interaction, and (2) the diagrams with three two-photon interactions. One representative of each type is shown in Figs. 2c and  2d. The subleading order diagrams are negligible at $\beta\approx 1\%$, but make modest quantitiative corrections on top of the leading order diagrams at large OD with $\beta\approx 5\%$. As the diagrams feature a loop structure, we refer to these as loop-order diagrams and their explicit calculation is presented in our accompanying work [WDM2025].
@@
Connected third-order correlation function.—Having identified the leading-order contributions to connected three-photon scattering, we now translate this into an experimentally accessible quantity that reflects both non-Gaussian intensity and field correlation in the outgoing light. The connected third-order correlation function is defined as
@@
where $g^{(2)}(x_{1},x_{2})$ and $g^{(3)}(x_{1},x_{2},x_{3})$ are the normalized second- and third-order correlation functions, respectively. For non-vanishing power, $g_{c}^{(3)}(x_{1},x_{2},x_{3})$ vanishes when either one photon is far away from the other two, or when the intensity correlation is Gaussian [Lemonde2014PRA, Jachymski2016, Stiesdal2018]. Moreover, if the state of light has only Gaussian field correlation, both $g^{(3)}$ and $g^{(2)}$ can be determined by the second-order moment and mean of $\hat{a}(x)$ and $\hat{a}^{\dagger}(x)$ by Isserlis’s theorem [Cardin2024, Shi2018, Hackl2021]. In the weak driving condition $\mathcal{O}(P_{\rm in}/\Gamma_{\rm tot})=\mathcal{O}(\beta)$  [^1] Stricly, we require $\mathcal{O}(P_{\rm in}/\Gamma_{\rm tot})=\mathcal{O}(\beta^{2})$, but it is shown in the accompanying work [WDM2025] that this condition can be loosened to be $\mathcal{O}(P_{\rm in}/\Gamma_{\rm tot})=\mathcal{O}(\beta)$., $g^{(n)}(x_{1},\dots,x_{n})=\expectationvalue{:\Pi_{j=1}^{n}\hat{a}^{\dagger}(x_{j})\hat{a}(x_{j}):}/\Pi_{k=1}^{n}\expectationvalue{\hat{a}^{\dagger}(x_{k})\hat{a}(x_{k})}$ for $n=2,3$ can be approximated as:
@@
where $\psi_{2}(x_{1},x_{2})=t_{0}^{2M}+\phi_{2}(x_{1},x_{2})$ with $\phi_{2}(x_{i},x_{j})$ is the entangled wavefunction generated by all possible two-photon interactions. The output power $\expectationvalue{\hat{a}^{\dagger}(x)\hat{a}(x)}\approx P_{\rm in}t_{0}^{2M}$, $\psi_{3}(x_{1},x_{2},x_{3})=t_{0}^{3M}+t_{0}^{M}\sum_{1\leq i<j\leq 3}\,\phi_{2}(x_{i},x_{j})+\phi_{3}(x_{1},x_{2},x_{3})$, where $\phi_{3}$ encodes the correlation induced by leading- and subleading-order three-photon interactions depicted in Fig. 2. We evaluate $\phi_{2}$ and $\phi_{3}$ analytically, but for brevity here, their expressions are given in [WDM2025].